March 27, 2013 Tracking CDC/FDC

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

Wednesday March 27, 2013 at 11:00am At Jefferson Lab, the meeting will be held in F326


To connect from the outside, please use ESNET

  1. ) ESNET: 8542553
  2. )To connect by telephone, dial:
  • You can look up toll-free number at
  • US and Canada: (866)740-1260 (toll free)
  • International: (303)248-0285 (toll call)
  • enter access code followed by the # sign: 3421244#

Hardware Agenda

  1. Announcement
  2. FDC Update
  3. CDC Update
  4. No Cost Extension to Contract??
    1. I have received emails on the contract coming to an end?
  5. Installation Details
    1. Based on the last installation meeting, it became clear that the FDC installation could move forward to parallel the electronics installation. Discussion of this point prior to the Friday meeting.


Present: Curtis, Naomi, Beni, Simon, Eugene

  1. FDC Update
    1. All four FDC packages are now tested. The last package (#2) was tested to satisfaction in the cosmic ray test stand. All strips and wires are operational. Package #3 in plane 5 has one HV sector that keeps tripping. The wire frame was taken out of the stack and put for tests between lexan sheets to isolate the problem. Some wires were found with some sort of coating on them towards the center of the frame. Removing these wires from the HV did not solve the trip problem. If anything it became worth. In a next step all wires of that sector will be de-soldered to check if the HV trips are related to the wires or to the board itself.
    2. Preparations are made to test the pre-amplifier cooling system in 126 using the new chiller operating with Fluorinert.
  2. fadc125
    1. Beni reported that JLab have received some fadc125 - testing is proceeding - many show problems with production - too much solder creating shorts etc - Beni will try to track down 3 for CMU :-) and talk to Cody about the updated fadc125lib. The first articles were delayed by 1 week, the real production will be 1 month late.
  3. CDC Update
    1. The VME PS returned earlier this week, Naomi hooked up 1 board on the CDC this morning - woohoo! There is a fair amount of pickup on the HVB alone, disconnected, providing its ground connection makes no difference to this, but the CDC signals look nice (apart from being an unexpected polarity but that is probably because of using the new cables with the old shaper unit).
    2. First signals!

      Magenta and green traces are CDC to old shaper unit to scope
      Straw A42 as seen by fadc125

  4. No Cost Extension to Contract
    1. Curtis received an email from JLab saying that the contract ends 4/30, Beni will communicate with JLab purchasing contract people directly.
  5. Installation Details
    1. The CDC team were hoping to deliver the CDC before the winter in order to avoid the bad weather during transit and have several months of testing in place with cosmics. The FDC installation is scheduled to be completed in Jan 2014, we were hoping that this could be brought forward.
    2. The FDC needs electronics in place to test it before its insertion into the magnet. Fernando had said (at the last installation meeting) that this would be available earlier than expected. Beni said that they would need at least 6 crates to test 1 package, can test 1 package at a time, then by the time they finished all the electronics would be available; the idea of inserting the FDC and then having to remove it, re-test and re-insert is very unwelcome. There is not room for the CDC on the platform until the FDC is in place and checked out. The FDC is scheduled to be moved onto the platform in August.
    3. The electronics should be ready by the end 2013 except for cpus, only 1/3 of these are arriving in FY13 due to budget issues (Eugene).
    4. Cabling trays are a bottleneck with manpower & space, routing of the cable trays more complicated than thought. Students could help with the cable bundling during the summer.
    5. The magnet delays are also preventing work from being done in the hall - magnet testing should be complete by June.
    6. Crates are to be filled elsewhere and tested before being brought into the hall.
    7. Curtis will remind Fernando & Matt about the next installation meeting.
    8. Eugene will email the latest schedule (or link to it online).